how do i find the equation of the line that passes through the points (-2,2) and (0,5)?
Answer by actuary(112)

Since two points are given, you first need to find the slope (m) of the line.
This is the general formula for the slope.
m=(y2-y1)/(x2-x1) Substituting the actual values into the formula we have
m = (5-2)/(0-(-2))= 3/2
Now you can use the point-slope formula
(y-2)=3/2*(x-(-2))
So y = 3/2x+3+2= (3/2)x+5
As a check, substitute the value for x=0 for the other given point.
y=(3/2)*0+5=5 (Checks out OK)
I hope that this helps